OpenStudy (anonymous):

Given the equation of the line (y-2)=3(x+1) what is the slope

OpenStudy (anonymous):

Can you give me the x and y ? insead of equation?

OpenStudy (anonymous):

It doesn't give a number for the x and y. It's Point-Slope Formula.

OpenStudy (anonymous):

y - 2 = 3 (x + 1) y - 2 - 1 = 3(x + 1 - 1) y - 1 = 3(x) y - 1 + 1 = 3 + 1(x) y = 4(x) Thats as far as I can get.

OpenStudy (anonymous):

If that is right let me know.

OpenStudy (anonymous):

\[y-2 = 3(x+1)\]\[y-2=3x+3\]\[y=3x+5\]slope = 3 over 1 or 3